OpenStudy (anonymous):

slope -6 through (-7,7) the equation of the line is?

Nnesha (nnesha):

slope intercept form is y = mx +b where m is slope and b is y-intercept so replace m by -6 and x y by order pair (x,y) then solve for b

OpenStudy (anonymous):

You can use point slope form y-y1=m(x-x1) to solve this as well

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ -7}}\quad ,&{\color{blue}{ 7}})\quad % (c,d) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= -6 \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

would this be (-6,0)

OpenStudy (anonymous):

what would be the equation of the line?

OpenStudy (anonymous):

y=6x+14 I think

Nnesha (nnesha):

(-6 , 0) is what ?? equation should be in this form y = mx +b

OpenStudy (anonymous):

y=-6x+0

OpenStudy (anonymous):

it is y=-6x-35

Nnesha (nnesha):

 replace x y and m by their values  so yes that's right :)